# Pathwise Test-Time Correction for Autoregressive Long Video Generation

### v1

@@ 5 Experiments
Baseline. We evaluate our test-time correction method on two baseline models, CausVid (Yin et al., 2025a) and Self-Forcing (Huang et al., 2025c). Both baselines are built on the Wan2.1-T2V-1.3B model (Wan et al., 2025) and generate 5-second video clips at 16 FPS a resolution of $832\times 480$.
Standard Evaluation. We benchmark our method against representative autoregressive video diffusion baselines, including CausVid (Yin et al., 2025a), Self-Forcing (Huang et al., 2025c), Rolling Forcing (Liu et al., 2025b), and LongLive (Yang et al., 2025). Following standard protocols, evaluations are conducted using VBench (Huang et al., 2024) on 128 prompts randomly sampled from MovieGen (Polyak et al., 2024). Unless otherwise specified, we conduct all experiments in the 30-second video generation setting, using Self-Forcing as the default baseline. More experimental settings and implementation details for fair comparison are included in the supplementary material.
Additional Evaluation. To rigorously evaluate temporal coherence, we complement standard VBench quality metrics with temporal color histograms and JEPA scores (Balestriero et al., 2025) to assess long-term temporal drift. Since temporal consistency scores can be artificially improved by suppressing motion, allowing models to effectively cheat the evaluation, we conduct comparisons under matched dynamic degrees to ensure a fair and meaningful assessment. In addition, we use t-LPIPS (Zhang et al., 2018) to explicitly measure visual discontinuities, serving as a direct proxy for flickering artifacts at autoregressive chunk boundaries in our ablation studies. Comprehensive experimental settings and implementation details for fair comparison are included in the supplementary material.
Qualitative Results. As shown in Figure 7, integrating our method with Self-Forcing and CausVid substantially reduces error accumulation in long-horizon video generation. While the original baselines exhibit temporal drift and visual degradation over time, our integration maintains stable temporal coherence and visual fidelity over 30-second sequences, especially in videos with complex motion and appearance changes. Under the 30-second setting, our training-free approach achieves visual quality comparable to, and in some cases better than, Rolling Forcing and LongLive, which rely on additional training or specialized mechanisms. These results demonstrate that our method provides an effective and general test-time solution for improving long-term temporal consistency in autoregressive video generation.
@@
Total NFE Timesteps VBench Metrics Boundary
750 500 250 Sub. Cons. Bg. Cons. Dyn. Deg. Mot. Sm. Img. Qual. Aes. Qual. t-LPIPS
@@
Quantitative Results. All quantitative results are summarized in Table 1. Under the 30-second generation setting on VBench, integrating our method into standard autoregressive baselines, Self-Forcing and CausVid, consistently improves long-horizon video generation quality across diverse prompts and scenes. In particular, our method substantially reduces error accumulation and temporal drift, leading to improved subject and background consistency while notably enhancing dynamic degree without sacrificing motion smoothness or imaging quality. Moreover, the proposed path-wise correction effectively stabilizes appearance evolution over time, as evidenced by lower color-shift L1 distances and higher histogram correlations between the first and last frames. At the semantic level, our method also improves JEPA consistency by reducing both the standard deviation and first–last score difference across the entire sequence, indicating more coherent long-term representations. Compared with training-based methods such as Rolling Forcing and LongLive, our approach achieves comparable long-horizon consistency and visual quality while preserving stronger motion dynamics and requiring no additional training or parameter updates at test time.
@@
Ablation Study on Path-wise Correction. We compare single-point and path-wise correction to evaluate the role of the stochastic sampling trajectory in practice. Single-point correction directly replaces the latent at a fixed denoising step, whereas path-wise correction re-noises the corrected prediction to the same noise level and resumes denoising along the original trajectory. As shown in Figure 9 and Table 4, single-point correction frequently introduces flickering and temporal instability, leading to degraded consistency metrics and higher t-LPIPS scores. In contrast, path-wise correction achieves consistently higher temporal consistency and substantially lower t-LPIPS, resulting in more stable videos with improved temporal coherence. These results demonstrate that effective test-time intervention requires integrating corrections along the sampling path rather than directly replacing latents.
Ablation Study on Noise-correction Steps. After establishing the necessity of path-wise correction, we evaluate how different numbers and placements of correction steps along the sampling path affect long-horizon generation quality. Enabling correction at noise levels 750, 500, and 250, either individually or in combination, consistently outperforms the baseline without correction, demonstrating the robustness and effectiveness of our method across different configurations. Considering the trade-off between performance and inference cost, we adopt correction at noise levels 500 and 250 in our experiments.
@@
Table 4: Comparison of correction strategies. We evaluate quality using VBench metrics alongside the Boundary metric.
Method VBench Metrics Boundary
Sub. Cons. Bg. Cons. Dyn. Deg. Mot. Sm. Img. Qual. Aes. Qual. t-LPIPS
@@
Method VBench Metrics
Sub. Cons. Bg. Cons. Dyn. Deg. Mot. Sm. Img. Qual. Aes. Qual.
@@ 6 Conclusion
In this paper, we propose Test-Time Correction, a training-free test-time method for stabilizing distilled autoregressive diffusion models in long-horizon video generation. The proposed approach addresses error accumulation by introducing training-free, reference-based correction along the stochastic sampling process, allowing corrected predictions to be smoothly inherited by subsequent denoising steps. Without modifying model parameters or requiring additional training, our method effectively suppresses temporal drift while preserving the original generation behavior. Extensive experiments demonstrate that Test-Time Correction consistently improves long-horizon stability across multiple distilled video generation models, extending the achievable generation length to 30 seconds with negligible computational overhead and competitive visual quality.
Impact Statement
This paper presents a training-free test-time method for improving the stability of autoregressive video generation models. The primary goal of this work is to advance the field of machine learning by enabling more reliable long-horizon video synthesis without additional training or model modification.
While the proposed method may contribute to downstream applications that rely on long video generation, such as content creation and simulation, it does not introduce new capabilities beyond existing video generative models. The potential societal impacts of this work are therefore aligned with those already associated with video generation technologies, and no specific additional ethical concerns are introduced by the method itself.
Appendix A Details on Samplers.
@@
Appendix B Details on Evaluations
Boundary Continuity (t-LPIPS). To measure perceptual discontinuities at segment junctions in autoregressive generation, we compute LPIPS (Zhang et al., 2018) only on boundary-adjacent frame pairs. Our autoregressive model generates a video as $K$ consecutive chunks. Let $t_{k}$ denote the last frame index of the $k$-th chunk; then the $k$-th boundary is the adjacent pair $(f_{t_{k}},f_{t_{k}+1})$ for $k\in\{1,\ldots,K-1\}$. We define the boundary score as the mean LPIPS over these K1K-1 pairs:
@@ Appendix B Details on Evaluations
Figure 10: JEPA-score and JEPA consistency for long-video evaluation.
@@
Appendix C Details on Methods.
Details on Test-time Optimization (TTO). Following the test-time adaptation protocols established in HyperNoise (Eyring et al., 2025) and AutoRefiner (Yu et al., 2025b), we perform gradient-based optimization at each sampling step. We employ an AdamW optimizer with a learning rate of $1\times 10^{-4}$. Specifically, at each denoising step for each latent chunk, the latent prediction is first decoded into pixel space via a pre-trained VAE decoder (Wan et al., 2025). We then compute the Mean Squared Error (MSE) loss and the CLIP score (Radford et al., 2021) on the decoded image with the initial image, which serve as proxies for pixel-level and semantic-level rewards, respectively, to guide the optimization process.
Appendix D Further Quantitative Results.
@@
Rolling Forcing (Liu et al., 2025b) 0.2956 0.5738 16.2365
Longlive (Yang et al., 2025) 0.3056 0.5969 16.8669
Self-Forcing (Huang et al., 2025d) 0.3548 0.5377 15.6360
@@
Dynamic Preservation. Adhering to the evaluation protocols established in AutoRefiner (Yu et al., 2025b), we conduct a comparative analysis of the dynamic degree against baseline methods, including Self-Forcing, Rolling Forcing, and Longlive. To rigorously assess the dynamic degree, we quantify the perceptual variation between temporally strided frames utilizing metrics such as LPIPS, SSIM, and PSNR with a fixed sampling interval $k$ (e.g., $k=12$). We model the magnitude of motion and structural evolution over time by computing the average distance $D=\mathbb{E}_{t}[\mathcal{M}(f_{t},f_{t+k})]$, where $\mathcal{M}$ represents the specific metric function (e.g., LPIPS) and $f_{t}$ denotes the frame at time step $t$. As evidenced in Tab. 7, unlike baseline approaches that often compromise motion magnitude to ensure stability, our method sustains a superior dynamic degree while preserving temporal coherence, thereby effectively maintaining the vividness of the generated content.
Appendix E Further Qualitative Results.

### v2

@@ 5 Experiments
5.1 Settings
Baseline. We evaluate our test-time correction method on two baseline models, CausVid [yin2025slow] and Self-Forcing [huang2025self]. Both baselines are built upon the Wan2.1-T2V-1.3B model [wan2025] and generate 5-second video clips at 16 FPS with a resolution of $832\times 480$. More implementation details for fair comparison are included in the supplementary material.
Standard Evaluation. We benchmark our method against representative autoregressive video diffusion baselines, including CausVid [yin2025slow], Self-Forcing [huang2025self], Rolling Forcing [liu2025rolling], and LongLive [yang2025longlive]. Following standard protocols, evaluations are conducted using VBench [huang2024vbench] on 128 prompts randomly sampled from MovieGen [polyak2024movie]. Unless otherwise specified, we conduct all experiments in the 30-second video generation setting, using Self-Forcing as the default baseline. More experimental settings and implementation details for fair comparison are included in the supplementary material.
Additional Evaluation. To rigorously evaluate temporal coherence, we complement standard VBench quality metrics with temporal color histograms and JEPA scores [JEPA] to assess long-term temporal drift. Since temporal consistency scores can be artificially improved by suppressing motion, allowing models to effectively cheat the evaluation, we conduct comparisons under matched dynamic degrees to ensure a fair and meaningful assessment. In addition, we use t-LPIPS [lpips] to explicitly measure visual discontinuities, serving as a direct proxy for flickering artifacts at autoregressive chunk boundaries in our ablation studies. Comprehensive experimental settings and implementation details for fair comparison are included in the supplementary material.
5.2 Results
Qualitative Results. As shown in Figure 7, integrating our method with Self-Forcing substantially reduces error accumulation in long-horizon video generation. While the original baselines exhibit temporal drift and visual degradation over time, our integration maintains stable temporal coherence and visual fidelity over 30-second sequences, especially in videos with complex motion and appearance changes. Under the 30-second setting, our training-free approach achieves visual quality comparable to, and in some cases better than, Rolling Forcing and LongLive, which rely on additional training or specialized mechanisms. These results demonstrate that our method provides an effective and general test-time solution for improving long-term temporal consistency in autoregressive video generation. More experimental settings and results are included in the supplementary material.
@@ 5.2 Results
Total NFE Timesteps VBench Metrics $\uparrow$ Boundary $\downarrow$
750 500 250 Subject Consistency Background Consistency Dynamic Degree Motion Smoothness Imaging Quality Aesthetic Quality t-LPIPS
@@
Quantitative Results. All quantitative results are summarized in Table 1 and Table 2. Under the 30-second generation setting on VBench, integrating our method into standard autoregressive baselines, Self-Forcing and CausVid, consistently improves long-horizon video generation quality across diverse prompts and scenes. In particular, our method substantially reduces error accumulation and temporal drift, leading to improved subject and background consistency while notably enhancing dynamic degree without sacrificing motion smoothness or imaging quality. Moreover, the proposed path-wise correction effectively stabilizes appearance evolution over time, as evidenced by lower color-shift L1 distances and higher histogram correlations between the first and last frames. At the semantic level, our method also improves JEPA consistency by reducing both the standard deviation and first–last score difference across the entire sequence, indicating more coherent long-term representations. Compared with training-based methods such as Rolling Forcing and LongLive, our approach achieves comparable long-horizon consistency and visual quality while preserving stronger motion dynamics and requiring no additional training or parameter updates at test time.
@@
Ablation Study on Pathwise Correction. We compare single-point and path-wise correction to evaluate the role of the stochastic sampling trajectory in practice. Single-point correction directly replaces the latent at a fixed denoising step, whereas path-wise correction re-noises the corrected prediction to the same noise level and resumes denoising along the original trajectory. As shown in Figure 9 and Table 5, single-point correction frequently introduces flickering and temporal instability, leading to degraded consistency metrics and higher t-LPIPS scores. In contrast, path-wise correction achieves consistently higher temporal consistency and substantially lower t-LPIPS, resulting in more stable videos with improved temporal coherence. These results demonstrate that effective test-time intervention requires integrating corrections along the sampling path rather than directly replacing latents.
Ablation Study on Noise-correction Steps. After establishing the necessity of path-wise correction, we evaluate how different numbers and placements of correction steps along the sampling path affect long-horizon generation quality as shown in Table 4. Enabling correction at noise levels 750, 500, and 250, either individually or in combination, consistently outperforms the baseline without correction, demonstrating the robustness and effectiveness of our method across different configurations. Balancing generation quality and inference overhead, we choose correction at noise levels 500 and 250 as our default, since this configuration yields consistent improvements in both quantitative metrics and qualitative visual fidelity.
@@
Table 5: Comparison of correction strategies, including single-point correction and pathwise correction. We evaluate quality using VBench metrics alongside the Boundary metric.
Method VBench Metrics $\uparrow$ Boundary $\downarrow$
Subject Consistency Background Consistency Dynamic Degree Motion Smoothness Imaging Quality Aesthetic Quality t-LPIPS
@@
Method VBench Metrics $\uparrow$
Subject Consistency Background Consistency Dynamic Degree Motion Smoothness Imaging Quality Aesthetic Quality
@@ 6 Conclusion
In this paper, we propose Test-Time Correction, a training-free test-time method for stabilizing distilled autoregressive diffusion models in long-horizon video generation. The proposed approach addresses error accumulation by introducing training-free, reference-based correction along the stochastic sampling process, allowing corrected predictions to be smoothly inherited by subsequent denoising steps. Without modifying model parameters or requiring additional training, our method effectively suppresses temporal drift while preserving the original generation behavior. Extensive experiments demonstrate that Test-Time Correction consistently improves long-horizon stability across multiple distilled video generation models, extending the achievable generation length to 30 seconds with a slight computational overhead and competitive visual quality.
Appendix 0.A Details on Samplers.
@@
Appendix 0.B Details on Evaluations
Boundary Continuity (t-LPIPS). To measure perceptual discontinuities at segment junctions in autoregressive generation, we compute LPIPS only on boundary-adjacent frame pairs. Our autoregressive model generates a video as $K$ consecutive chunks. Let $t_{k}$ denote the last frame index of the $k$-th chunk; then the $k$-th boundary is the adjacent pair $(f_{t_{k}},f_{t_{k}+1})$ for $k\in\{1,\ldots,K-1\}$. We define the boundary score as the mean LPIPS over these K1K-1 pairs:
@@ Appendix 0.B Details on Evaluations
Figure 10: JEPA-based long-video evaluation pseudocode. Given frames $I_{1:T}$ and a frozen encoder $\phi(\cdot)$, we compute JEPA-score from $\mathrm{svd}(\nabla_{X}\phi(X))$ and temporal drift statistics from first-frame-anchored feature cosine distances.
@@
Appendix 0.C Details on Methods.
Details on Test-time Optimization (TTO). Following the test-time adaptation protocols established in HyperNoise and AutoRefiner, we perform gradient-based optimization at each sampling step. We employ an AdamW optimizer with a learning rate of $1\times 10^{-4}$. Specifically, at each denoising step for each latent chunk, the latent prediction is first decoded into pixel space via a pre-trained VAE decoder. We then compute the Mean Squared Error (MSE) loss and the CLIP score on the decoded image with the initial image, which serve as proxies for pixel-level and semantic-level rewards, respectively, to guide the optimization process.
Appendix 0.D Further Quantitative Results.
@@
Rolling Forcing 0.2956 0.5738 16.2365
Longlive 0.3056 0.5969 16.8669
Self-Forcing 0.3548 0.5377 15.6360
@@
Dynamic Preservation. Adhering to the evaluation protocols established in AutoRefiner, we conduct a comparative analysis of the dynamic degree against baseline methods, including Self-Forcing, Rolling Forcing, and Longlive. To rigorously assess the dynamic degree, we quantify the perceptual variation between temporally strided frames utilizing metrics such as LPIPS, SSIM, and PSNR with a fixed sampling interval $k$ (e.g., $k=12$). We model the magnitude of motion and structural evolution over time by computing the average distance $D=\mathbb{E}_{t}[\mathcal{M}(f_{t},f_{t+k})]$, where $\mathcal{M}$ represents the specific metric function (e.g., LPIPS) and $f_{t}$ denotes the frame at time step $t$. As evidenced in Tab. 8, unlike baseline approaches that often compromise motion magnitude to ensure stability, our method sustains a superior dynamic degree while preserving temporal coherence, thereby effectively maintaining the vividness of the generated content.
Appendix 0.E Further Qualitative Results.